autonomous consumption c1=R2000
autonomous investment =R6000
marginal propensity to consume c=3/4
the value of the multiplier is
(1) 3/4
(2) 1/4
(3) 5
(4) 4
1
Expert's answer
2019-09-16T09:31:46-0400

(4) 4

If MPC = 3/4 or 0.75, multiplier is

=1/(1-0.75) = 1/0.25 = 100/25 = 4